# Bringing Back the Context: Camera Trap Species Identification as Link Prediction on Multimodal Knowledge Graphs

### v3

@@ 3.3 Model Architecture
3.3.1 Multi-modality Encoders
We use an ImageNet pre-trained ResNet-50 [24] as the image encoder. The base feature of each location is represented as a 2D vector [latitude, longitude]. Following prior work [49], we use an MLP to project the 2D location feature to a higher dimensional space. Similarly, for temporal context, we use an MLP to project the integer value of the hour timestamp to the higher dimensional embedding space. For categorical entities such as species labels and taxa, we learn dense embeddings as representations.
3.3.2 Training
We train the model using an optimization strategy based on the modality of the tail entity. For categorical attributes, we formulate it as a multi-class classification problem and use standard cross-entropy loss to train the model. For instance, in case of a given image-species label ground truth triple (,instance $(\mathcal{I},\texttt{instance\;of},s)$, the loss is defined as:
(,instance $\begin{split}\mathcal{L}(\mathcal{I},&\texttt{instance\;of},s)=\\ &-\log\frac{\exp(\psi(\mathcal{I},\texttt{instance\;of},s))}{\sum_{s^{{}^{\prime}}\in S}\exp(\psi(\mathcal{I},\texttt{instance\;of},s^{{}^{\prime}}))},\end{split}$ (2)
@@ 3.3.2 Training
For numerical attributes such as location and time, we formulate it as a multi-class multi-label classification problem and use a binary cross-entropy loss to optimize the parameters. This choice is motivated by the fact that images can be associated with a range of GPS coordinates and timestamps, e.g., most animals are active multiple times during the day. The label space comprises all entities of ground truth modality. For instance, in the case of a given time modality ground truth triple $(\mathcal{I},\texttt{time},t)$, the loss is defined as:
$\begin{split}\mathcal{L}(\mathcal{I},\texttt{time},t)=-\sum_{t^{{}^{\prime}}}l^{\mathcal{I},time}_{t^{{}^{\prime}}}\cdot\log(\sigma(\psi(\mathcal{I},\texttt{time},t^{{}^{\prime}})))+\\ (1-l^{\mathcal{I},time}_{t^{{}^{\prime}}})\cdot(1-\log(\sigma(\psi(\mathcal{I},\texttt{time},t^{{}^{\prime}})))),\end{split}$ (3)
where $l^{\mathcal{I},time}_{t^{{}^{\prime}}}$ is a binary label that indicates whether the triple $(\mathcal{I},time,t^{{}^{\prime}})$ exists in the set of observed triples and $\sigma(\cdot)$ is the sigmoid activation function. We train the model by sequentially minimizing the objective on each type of context triple. Figure 1 illustrates the overall model architecture.
4 Experimental Setup
4.1 Datasets
We test our approach on the iWildCam2020-WILDS dataset [30], a variant of the iWildCam 2020 dataset [9] and Snapshot Mountain Zebra [48]. iWildCam2020-WILDS is a benchmark dataset designed to test out-of-distribution (OOD) generalization for the task of species classification. The label space consists of 182 species. Each domain corresponds to a different location of the camera trap. The training and test images belong to disjoint sets of locations in the OOD setting.
Model Multi-modality Val. Acc. Test Acc.
@@ 4.1 Datasets
Empirical Risk Minimization (ERM) [30] 62.7 (2.4) 71.6 (2.5)
CORAL [59] 60.3 (2.8) 73.3 (4.3)
Group DRO [26] 60.0 (0.7) 72.7 (2.0)
Fish [56] 58.0 (0.2) 63.2 (0.7)
ABSGD [50] 72.7 (1.8)
COSMO (no-context) 63.2 (0.4) 68.8 (2.1)
Single
context
COSMO 62.8 (2.2) (-0.4) 72.4 (2.5) (+3.6)
64.4 (1.0) (+1.2) 74.5 (3.6) (+5.7)
64.7 (0.4) (+1.5) 71.1 (3.1) (+2.3)
Multiple
contexts
COSMO 65.4 (0.4) (+2.2) 70.4 (2.1) (+1.6)
64.9 (1.6) (+1.7) 73.7 (3.8) (+4.9)
63.0 (2.1) (-0.2) 74.2 (2.2) (+5.4)
65.0 (1.6) (+1.8) 71.5 (2.8) (+2.7)
Table 1: Species Classification results on iWildCam2020-WILDS (OOD) dataset. The first baseline in the second section shows the no-context baseline that uses only image-species labels as KG edges. All models use a pre-trained ResNet-50 as image encoder. Parentheses show standard deviation across 3 random seeds. Missing values are denoted by –.
@@
The Snapshot Mountain Zebra dataset, part of the Snapshot Safari project [48], consists of camera trap images captured at the Mountain Zebra National Park, South Africa. The label space consists of 53 species, mostly annotated at the species level. Prominent animal species include Cape Mountain zebra, kudu, and springbok. We manually split the images to have disjoint camera traps in each split due to the absence of a standard split. These datasets pose a significant challenge for species recognition due to factors like inadequate illumination, motion blur, occlusion, temporal variations, and diverse weather conditions, effectively reflecting the complexities of real-life camera trap usage. Dataset statistics are shown in Table 2.
4.2 Baselines
@@
We use the COSMO with no context that uses just the species label edges as our baseline. In addition, we compare with the following baseline algorithms for OOD generalization: Empirical Risk Minimization (ERM) [30], which trains the model to minimize average training loss, CORAL [59], a method for unsupervised domain adaptation that learns domain invariant features, Group DRO [26], an algorithm that uses distributionally robust optimization to perform well on subpopulation shifts, Fish [56] that attempts domain adaptation using gradient matching, and ABSGD [50], an optimization method for addressing data imbalance. All models use a pre-trained ResNet-50 as image encoder. We evaluate the models using overall accuracy as the metric.
4.3 Implementation Details
@@
We implement our models in PyTorch. The hidden dimension of the multimodal KG embedding model is set to 512. The images are resized to $448\times 448$ before input to the image encoder. We use Adam [29] optimizer with a learning rate of 3e-5 and 1e-3 for the image encoder and the rest of the parameters, respectively. We use early stopping based on validation accuracy to prevent overfitting. All results are reported with averages across three random seeds. A more comprehensive description of the hyperparameter setup is given in the Appendix.
@@
Table 2: Dataset Statistics.
Model Multi-modality Test Acc.
@@
ERM [30] 96.2 (0.6)
CORAL [59] 96.6 (1.2)
Group DRO [26] 93.4 (2.1)
ABSGD [50] 93.4 (2.0)
COSMO (no-context) 92.9 (2.5)
COSMO 93.9 (2.8) (+1.0)
95.3 (3.1) (+2.4)
96.8 (0.4) (+3.9)
Table 3: Species Classification results on Snapshot Mountain Zebra dataset. We obtain the results for OOD baselines by training them on this dataset using pubicly available code.
5 Results
@@
Figure 2: Comparison of COSMO model with and without taxonomy edges (iWildCam2020-WILDS validation set). The use of taxonomy information helps the model to avoid semantically implausible predictions.
5.1 Performance Comparison with Addition of Multimodal Context
We add taxonomy, location, and temporal context information to the base KG and observe the impact on the species classification performance. Table 1 shows the results for iWildCam2020-WILDS dataset. We make the following observations from these results:
@@ 5.1 Performance Comparison with Addition of Multimodal Context
Secondly, we observe that the use of multiple contexts results in a performance boost in a majority of cases. For instance, the addition of location and time attributes improves over the taxonomy baseline by a margin of $2.6$% and $2.1$% respectively in terms of validation set accuracy. Similarly, the taxonomy with time baseline obtains an improvement of $1.3$% and $2.6$% over the taxonomy and time baselines, respectively in terms of test accuracy.
@@
5.2 Comparison with OOD Generalization Approaches
We compare the performance of the COSMO with methods specifically designed for out-of-domain generalization. Notably, our best-performing model, which uses location as context, achieves state-of-the-art performance in terms of OOD test accuracy, outperforming the existing SOTA model by $1.2$% on the iWildCam2020-WILDS dataset. Likewise, COSMO with taxonomy and time contexts outperforms existing approaches on Snapshot Mountain Zebra dataset. This demonstrates the effectiveness of leveraging diverse multimodal contexts for achieving more robust OOD generalization, even in the absence of sophisticated objectives aimed at improving domain generalization, e.g., CORAL, Group DRO, ABSGD, and Fish.
@@ 5.2 Comparison with OOD Generalization Approaches
Table 4: Performance comparison among different KGE models (iWildCam2020-WILDS). DistMult outperforms ConvE in a majority of cases.
@@
Table 5: Quantitative evaluation of COSMO errors with and without taxonomy using a hierarchical distance metric (iWildCam2020-WILDS). The taxonomy-aware model achieves better performance in terms of Avg. LCA height.
5.3 Compairson with Alternative KGE backbones
@@
In our preliminary experiments, we also experiment with ConvE [17], a strong neural network baseline, as an alternate choice for the KGE backbone model, instead of DistMult for the iWildCam2020-WILDS dataset. We observe that DistMult outperforms ConvE in a majority of cases, especially when all context types are used together. Furthermore, DistMult holds the advantage of being more computationally efficient than neural network based KG embedding approaches.
5.4 Fine-grained Analyses
5.4.1 Error Analysis for the Taxonomy-aware Model
@@
To analyze the predictions of our model with and without taxonomy information, we employ a metric that takes into account the hierarchical structure of the labels. Conventional measures like top-1 accuracy treat all errors equally, disregarding the semantic relationships among labels. Hence, we use the Least Common Ancestor (LCA) [10] for the misclassified examples as the metric for this analysis (Table 5). A lower LCA value indicates that the errors made by the taxonomy-aware model are more semantically related to the true label than the baseline.
@@
We compare the predictions of COSMO which uses taxonomy to the no-context baseline (Figure 2). Notably, the inclusion of taxonomy information assists the model in avoiding implausible predictions. For instance, consider the case of the animal ocelot (eopardus pardalis), which belongs to the cat family (feliformia). The use of taxonomy information prevents the misprediction of this animal as a gray fox, which belongs to the dog family (caniformia). Similarly, in the second example, the baseline model incorrectly predicts the given image as Central American agouti, a mammal, instead of ocellated turkey, a bird.
@@
(a) Each color square shows the distance between the corresponding validation cluster centroid on x-axis and the training cluster centroid on y-axis. The correlation peaks along the diagonal [^3] The null value in row 4 is due to the absence of species overlap with respective validation clusters. The null value in columns 3 and 4 indicates the absence of these clusters in the validation set..
@@
(b) Each color square shows the distance between the corresponding training hour slot on x-axis and validation hour slot on y-axis. The correlation peaks for day-day and night-night hour slots.
@@
Figure 3: Correlation analysis for location and time attributes. Best viewed in color.
@@
Figure 4: Plot of location GPS coordinates for training and validation splits (iWildCam2020-WILDS). The coordinates can be grouped into six clusters. Most coordinates exhibit an overlap with their respective cluster centroids at this visualization scale.
5.4.2 Correlation Analysis for Location and Time Attributes
@@
We examined the relationship between species distribution and numerical attributes, such as location and time, to gain insights into how these contexts contribute to the task. The location coordinates can be grouped into six clusters. A visualization of the location clusters is shown in Figure 4. For each pair of cluster centroids, we compute the Bhattacharyya distance [11], a measure of similarity between probability distributions, between the training and validation set species distributions (Figure 3). Similarly, we plot the distance between species distributions corresponding to each hour of the day (Figure 2(b)). We observe that the similarity (corresponds to lower distance) peaks along the diagonal for the location attribute, as well as for the day/night categorization of the time attribute. This suggests these metadata give a prior for species class distribution.
@@
Table 6: Performance comparison on under-represented species classification (OOD Test set). The best performing COSMO model improves over the ERM baseline by a significant margin.
5.4.3 Performance Comparison for Under-represented Species

### v5

@@ 3.3. Model Architecture
3.3.1. Multi-modality Encoders
We use an ImageNet pre-trained ResNet-50 (He et al., 2016) as the image encoder. The base feature of each location is represented as a 2D vector [latitude, longitude]. Following prior work (Pezeshkpour et al., 2018), we use an MLP to project the 2D location feature to a higher dimensional space. Similarly, for temporal context, we use an MLP to project the integer value of the hour timestamp to the higher dimensional embedding space. For categorical entities such as species labels and taxa, we learn dense embeddings as representations.
3.3.2. Training
We train the model using an optimization strategy based on the modality of the tail entity. For categorical attributes, we formulate it as a multi-class classification problem and use standard cross-entropy loss to train the model. For instance, in case of a given image-species label ground truth triplet (,instance $(\mathcal{I},\texttt{instance\;of},s)$, the loss is defined as:
(2) (,instance $\begin{split}\mathcal{L}(\mathcal{I},&\texttt{instance\;of},s)=\\ &-\log\frac{\exp(\psi(\mathcal{I},\texttt{instance\;of},s))}{\sum_{s^{{}^{\prime}}\in S}\exp(\psi(\mathcal{I},\texttt{instance\;of},s^{{}^{\prime}}))},\end{split}$
@@ 3.3.2. Training
For numerical attributes such as location and time, we formulate it as a multi-class multi-label classification problem and use a binary cross-entropy loss to optimize the parameters. This choice is motivated by the fact that images can be associated with a range of GPS coordinates and timestamps, e.g., most animals are active multiple times during the day. The label space comprises all entities of ground truth modality. For instance, in the case of a given time modality ground truth triplet $(\mathcal{I},\texttt{time},t)$, the loss is defined as:
(3) $\begin{split}\mathcal{L}(\mathcal{I},\texttt{time},t)=-\sum_{t^{{}^{\prime}}}l^{\mathcal{I},time}_{t^{{}^{\prime}}}\cdot\log(\sigma(\psi(\mathcal{I},\texttt{time},t^{{}^{\prime}})))+\\ (1-l^{\mathcal{I},time}_{t^{{}^{\prime}}})\cdot(1-\log(\sigma(\psi(\mathcal{I},\texttt{time},t^{{}^{\prime}})))),\end{split}$
where $l^{\mathcal{I},time}_{t^{{}^{\prime}}}$ is a binary label that indicates whether the triplet $(\mathcal{I},time,t^{{}^{\prime}})$ exists in the set of observed triplets and $\sigma(\cdot)$ is the sigmoid activation function. We train the model by sequentially minimizing the objective on each type of context triplet. Figure 1 illustrates the overall model architecture.
4. Experimental Setup
4.1. Datasets
We test our approach on the iWildCam2020-WILDS dataset (Koh et al., 2021), a variant of the iWildCam 2020 dataset (Beery et al., 2020) and Snapshot Mountain Zebra (Pardo et al., 2021). iWildCam2020-WILDS is a benchmark dataset designed to test OOD generalization for the task of species classification. The label space consists of 182 species. Each domain corresponds to a different location of the camera trap. The training and test images belong to disjoint sets of locations in the OOD setting.
Table 1. Species classification results on iWildCam2020-WILDS (OOD) dataset. The first baseline in the second section shows the no-context baseline that uses only image-species labels as KG edges. All models use a pre-trained ResNet-50 as image encoder. Parentheses show standard deviation across 3 random seeds. We highlight the best result in bold and the second best with underline. We mark the improvements over COSMO (no-context) in green. Missing values are denoted by –.
Model Multi-modality Val. Acc. (%) Test Acc. (%)
@@ 4.1. Datasets
Empirical Risk Minimization (ERM) (Koh et al., 2021) 62.7 ($\pm$2.4) 71.6 ($\pm$2.5)
CORAL (Sun and Saenko, 2016) 60.3 ($\pm$2.8) 73.3 ($\pm$4.3)
Group DRO (Hu et al., 2018) 60.0 ($\pm$0.7) 72.7 ($\pm$2.0)
Fish (Shi et al., 2022) 58.0 ($\pm$0.2) 63.2 ($\pm$0.7)
ABSGD (Qi et al., 2023) 72.7 ($\pm$1.8)
MLP-concat 27.3 ($\pm$0.8) 39.6 ($\pm$1.0)
COSMO (no-context) 63.2 ($\pm$0.4) 68.8 ($\pm$2.1)
Single context COSMO 62.8 ($\pm$2.2) (-0.4) 72.4 ($\pm$2.5) (+3.6)
64.4 ($\pm$1.0) (+1.2) 74.5 ($\pm$3.6) (+5.7)
64.7 ($\pm$0.4) (+1.5) 71.1 ($\pm$3.1) (+2.3)
Multiple contexts COSMO 65.4 ($\pm$0.4) (+2.2) 70.4 ($\pm$2.1) (+1.6)
64.9 ($\pm$1.6) (+1.7) 73.7 ($\pm$3.8) (+4.9)
63.0 ($\pm$2.1) (-0.2) 74.2 ($\pm$2.2) (+5.4)
65.0 ($\pm$1.6) (+1.8) 71.5 ($\pm$2.8) (+2.7)
@@
Snapshot Mountain Zebra comprises camera trap images taken at the Mountain Zebra National Park in South Africa as a part of the Snapshot Safari project (Pardo et al., 2021). The label space consists of 53 species, mostly annotated at the species level. Prominent animal species include Cape Mountain zebra, kudu, and springbok. The location coordinates are not available for this dataset due to privacy and security reasons. We manually split the images to have disjoint camera traps in each split due to the absence of a standard split. These datasets pose a significant challenge for species recognition due to factors like inadequate illumination, motion blur, occlusion, temporal variations, and diverse weather conditions, effectively reflecting the complexities of real-life camera trap usage. Dataset statistics are shown in Table 2.
4.2. Baselines
@@
We use the COSMO with no context that uses just the species label edges as our baseline. In addition, we compare with the following baseline algorithms for OOD generalization: Empirical Risk Minimization (ERM) (Koh et al., 2021), which trains the model to minimize average training loss, CORAL (Sun and Saenko, 2016), a method for unsupervised domain adaptation that learns domain invariant features, Group DRO (Hu et al., 2018), an algorithm that uses distributionally robust optimization to perform well on subpopulation shifts, Fish (Shi et al., 2022) that attempts domain adaptation using gradient matching, and ABSGD (Qi et al., 2023), an optimization method for addressing data imbalance. As an alternative way of incorporating contextual information, we implement MLP-concat, a baseline which utilizes the location and temporal features at both training and inference time. It uses vanilla concatenation to fuse visual and spatiotemporal representations which are then fed into an MLP. The missing features are substituted by a mean value computed over the training dataset. All models use a pre-trained ResNet-50 as image encoder. We evaluate the models using overall accuracy as the metric.
4.3. Implementation Details
@@
We implement our models in PyTorch. The hidden dimension of the multimodal KG embedding model is set to 512, with a batch size of 16. The images are resized to $448\times 448$ before input to the image encoder. For the location and time attributes, we use a 3-layer MLP that projects the feature input dimension to the embedding dimension and uses PReLU as the activation function. We use Adam (Kingma and Ba, 2015) optimizer with a learning rate of 3e-5 and 1e-3 for the image encoder and the rest of the parameters, respectively. In our experiments, the models on iWildCam2020-WILDS and Snapshot Mountain Zebra were trained for 12 and 15 epochs, respectively. We use early stopping based on validation accuracy to prevent overfitting. The early stopping patience parameter is set to 5 epochs. All results are reported with averages across three random seeds.
Table 2. Dataset Statistics.
@@
Table 3. Species classification results on Snapshot Mountain Zebra dataset. We obtain the results for OOD baselines by training them on this dataset using publicly available code. We mark the improvements over COSMO (no-context) in green.
Model Multi-modality Test Acc. (%)
@@
ERM (Koh et al., 2021) 96.2 ($\pm$0.6)
CORAL (Sun and Saenko, 2016) 96.6 ($\pm$1.2)
Group DRO (Hu et al., 2018) 93.4 ($\pm$2.1)
ABSGD (Qi et al., 2023) 93.4 ($\pm$2.0)
MLP-concat 94.7 ($\pm$0.0)
COSMO (no-context) 92.9 ($\pm$2.5)
COSMO 93.9 ($\pm$2.8) (+1.0)
95.3 ($\pm$3.1) (+2.4)
96.8 ($\pm$0.4) (+3.9)
5. Results
@@
5.1. Performance Comparison with Addition of Multimodal Context
We add taxonomy, location, and temporal context information to the base KG and observe the impact on the species classification performance. Table 1 shows the results for the iWildCam2020-WILDS dataset. We make the following observations from these results:
@@ 5.1. Performance Comparison with Addition of Multimodal Context
Secondly, we observe that the use of multiple contexts results in a performance boost in a majority of cases. For instance, the addition of location and time attributes improves over the taxonomy baseline by a margin of $2.6$% and $2.1$% respectively in terms of the validation set accuracy. Similarly, the taxonomy with time baseline obtains an improvement of $1.3$% and $2.6$% over the taxonomy and time baselines, respectively in terms of test accuracy.
@@
5.2. Comparison with OOD Generalization Approaches
We compare the performance of the COSMO with methods specifically designed for out-of-domain generalization. Notably, our best-performing model, which uses location as context, achieves state-of-the-art performance in terms of OOD test accuracy, outperforming the existing SOTA model (CORAL) by $1.2$% on the iWildCam2020-WILDS dataset. Likewise, COSMO with taxonomy and time contexts outperforms existing approaches on the Snapshot Mountain Zebra dataset. This demonstrates the effectiveness of leveraging diverse multimodal contexts for achieving more robust OOD generalization, even in the absence of sophisticated objectives aimed at improving domain generalization, e.g., CORAL, Group DRO, ABSGD, and Fish. The MLP-concat baseline overfits the training camera trap locations on the iWildCam2020-WILDS dataset, resulting in suboptimal performance. COSMO consistently outperforms the MLP-concat baseline by a significant margin across both datasets.
Table 4. Performance comparison among different KGE models (iWildCam2020-WILDS). DistMult outperforms ConvE in a majority of cases.
@@ 5.2. Comparison with OOD Generalization Approaches
Figure 2. Comparison of COSMO model with and without taxonomy edges (iWildCam2020-WILDS validation set). The use of taxonomy information helps the model to avoid semantically implausible predictions.
Table 5. Quantitative evaluation of COSMO errors with and without taxonomy using a hierarchical distance metric (iWildCam2020-WILDS). The taxonomy-aware model achieves better performance in terms of Avg. LCA height.
@@
5.3. Compairson with Alternative KGE Backbones
@@
In our preliminary experiments, we explored the use of ConvE (Dettmers et al., 2018), a strong neural network baseline, as an alternative to DistMult for the KGE backbone model for the iWildCam2020-WILDS dataset (Table 4). Sun et al. (Sun et al., 2020) show that ConvE outperforms more recent neural network KGE models when evaluated properly. We observe that DistMult outperforms ConvE in a majority of cases, particularly when all incorporating all context types simultaneously. Furthermore, DistMult offers the advantage of being more computationally efficient than neural network based KG embedding approaches.
5.4. Fine-grained Analyses
5.4.1. Error Analysis for the Taxonomy-aware Model
@@
To analyze the predictions of our model with and without taxonomy information, we employ a metric that takes into account the hierarchical structure of the labels. Conventional measures like top-1 accuracy treat all errors equally, disregarding the semantic relationships among labels. Hence, we use the Least Common Ancestor (LCA) (Bertinetto et al., 2020) for the misclassified examples as the metric for this analysis (Table 5). A lower LCA value indicates that the errors made by the taxonomy-aware model are more semantically related to the true label compared to the baseline.
@@
We compare the predictions of COSMO which uses taxonomy to the no-context baseline (Figure 2). Notably, the inclusion of taxonomy information assists the model in avoiding implausible predictions. For instance, consider the case of the animal ocelot (Leopardus pardalis), which belongs to the cat family (feliformia). The use of taxonomy information prevents the misprediction of this animal as a gray fox, which belongs to the dog family (caniformia). Similarly, in the second example, the baseline model incorrectly predicts the given image as Central American agouti, a mammal, instead of ocellated turkey, a bird.
@@
(a) Each color square shows the distance between the corresponding validation cluster centroid on x-axis and the training cluster centroid on y-axis. The correlation peaks along the diagonal (highlighted in red) [^3] The null value in row 4 is due to the absence of species overlap with respective validation clusters. The null value in columns 3 and 4 indicates the absence of these clusters in the validation set..
@@
(b) Each color square shows the distance between the corresponding training hour slot on x-axis and validation hour slot on y-axis. The correlation peaks for day-day and night-night hour slots (highlighted in red).
@@
Figure 3. Correlation analysis for location and time attributes. Best viewed in color.
@@
Figure 4. Plot of location GPS coordinates for training and validation splits (iWildCam2020-WILDS). The coordinates can be grouped into six clusters. Most coordinates exhibit an overlap with their respective cluster centroids at this visualization scale. Best viewed in color.
5.4.2. Correlation Analysis for Location and Time Attributes
@@
We examined the relationship between species distribution and numerical attributes, such as location and time, to gain insights into how these contexts contribute to the task. The location coordinates can be grouped into six clusters. A visualization of the location clusters is shown in Figure 4. For each pair of cluster centroids, we compute the Bhattacharyya distance (Bhattacharyya, 1946), a measure of similarity between probability distributions, between the training and validation set species distributions (Figure 3). Similarly, we plot the distance between species distributions corresponding to each hour of the day (Figure 3(b)). We observe that the similarity (corresponds to lower distance) peaks along the diagonal for the location attribute, as well as for the day/night categorization of the time attribute. This suggests these metadata give a prior for species class distribution.
Table 6. Performance comparison on under-represented species classification (OOD test set). The best performing COSMO model improves over the ERM baseline by a significant margin.
@@
5.4.3. Performance Comparison for Under-represented Species